imagine a girl of mass 50kg standing on pencil heels,each of area of cross-section of 1cm2 and an elephant of mass 2000kg and foot area of 250cm2 standing on floor. find the pressure exerted by each of them assuming acceleration due to gravity is 10m/s2. Which of the two exerts more pressure on the floor?
1
Expert's answer
2018-07-16T15:29:08-0400
Wright of girl:
W_g=m_g∙g=50 kg∙10 m/s^2=500 N
Since girl has two feet, the total area of girl’s heels is:
S_g=2S_h=2 cm^2=0.0002 m^2
Thus the pressure exerted by girl is given by
P_g=W_g/S_g =(500 N)/(0.0002 m^2 )=25 bar
Weight of elephant:
W_e=M_e∙g=2000 kg∙10 m/s^2=20000 N
Total area of elephants feet (4 feet):
S_e=4S_f=4∙250 cm^2=1000 cm^2=0.1 m^2
Pressure exerted by elephant:
P_e=W_e/S_e =(20000 N)/(0.1 m^2 )=2 bar
Therefore the girl exerts more pressure on the floor than the elephant.
Answer: girl: 25 bar, elephant: 2 bar.